OpenStudy (anonymous):

A number divides 206, 368, and 449 leaving the same remainder in each cases.The largest such number is?

OpenStudy (anonymous):

the answer is 81

OpenStudy (apoorvk):

Sorry I made a mistake. :/ okay let the remainder be 'x'. Now that basically means that if the three nos. had been (206-x), (368-x) and (449-x), they would have been exactly divisible by the divisor

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

Small typo, (368 - 206, 368 - 449)=81

OpenStudy (anonymous):

@FoolForMath how u got this (368 - 206, 368 - 449)=81 can u explain

OpenStudy (anonymous):

can u explain @FoolForMath

OpenStudy (anonymous):

I surely can, but the proof is somewhat big, sorry.

OpenStudy (anonymous):

can any one solve this????

OpenStudy (anonymous):

wait for asnaseer the big man!

OpenStudy (anonymous):

@FoolForMath thanzz for ur effort to help me..

OpenStudy (asnaseer):

I think this is the method what FFM is employing: If 206 leaves some remainder after being divided by n, then we know: 206 - r = na where r is the remainder and 'a' is the multiple of n. therefore: 368 = 206 + 162 = na + 162 implies that 162 must also be exactly divisible by n. similarly: 449 = 368 + 81 = na + 162 + 81 implies 81 must also be exactly divisible by n

OpenStudy (asnaseer):

therefore, the largest such multiple must be 81

OpenStudy (asnaseer):

hope it makes sense shameer1

OpenStudy (anonymous):

thznzzz

OpenStudy (asnaseer):

yw - and thx to FFM for introducing me to yet another new concept :)